OpenStudy (anonymous):

Partial fractions 2x ------ (x+1)^3 I keep getting 2x = A(x+1)^2 + B(x+1) -2, but I can't solve for A and B. I do know that A+B = 2.

OpenStudy (anonymous):

2x/(x+1)^3=A/(x+1)+B/(x+1)^2+ C/(x+1)^3 2x = A(x+1)^2 + B(x+1) +C is correct. Now set up a system of equations based on degrees of x: x^2: 0=A x: 2=2A+B constants: 0=A+B+C Solving through substitution you get A=0, B=2, C=-2 So 2x/(x+1)^3=2/(x+1)^2 -2/(x+1)^3

OpenStudy (anonymous):

What was that system of equations you did? I'm not sure I understand your method.

OpenStudy (anonymous):

@bettirocked

OpenStudy (phi):

I keep getting 2x = A(x+1)^2 + B(x+1) -2 expand A(x+1)^2 to get Ax^2 +2Ax + A same for B: Bx+B we get 0x^2 + 2x + 0 = Ax^2 +2Ax+Bx + A+B-2 0x^2 + 2x + 0 = Ax^2 +(2A+B)x + (A+B-2) now equate the coefficients of each power of x term: 0 x^2 = A x^2 or 0 = A 2x = (2A+B)x or 2= 2A+B 0 = (A+B-2)

OpenStudy (anonymous):

Ohh.....I forgot to add 0x^2 to the left! Thanks Phi!